OpenStudy (anonymous):

the x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line?

OpenStudy (anonymous):

\[(-5;0) and (0;-2)\] \[m=\frac{ -2-0 }{ 0-(-5) }=\frac{ -2 }{ 5 }\] \[y=mx+c\] \[using the point (0;5) 0= \frac{ -2 }{ 5 }x+c then 2+c=0 \therefore c=-2\] \[\therefore y = \frac{ -2 }{ 5 }x-2\]